OpenStudy (anonymous):

given the equation 2x+4=4x-2 select the reasoning that correctly solves for x

OpenStudy (anonymous):

Select the reasoning? lol So we have option ( ^.^)

OpenStudy (anonymous):

4x-2x-2=4 4x-2x=4+2 2x=6 x=6/2 x=3

OpenStudy (anonymous):

add 2 subtract 2x then divide by 2 add 2 subtract 4x then divide by -2 subtract 4 subtract 2x then divide by -2 subtract 4 subtract 4x then divide by 2

OpenStudy (anonymous):

is it a?

OpenStudy (anonymous):

Do you believe it is A? did we add 2 to both sides? Then did we subtract 2x from both side? Then did we divide by 2 to get the answer of 3?

OpenStudy (anonymous):

yea i believe so

OpenStudy (anonymous):

Then your right.

OpenStudy (anonymous):

ok :) thnxs!! agian

OpenStudy (anonymous):

No problem.